OpenStudy (anonymous):

Solve the triangle. a = 12, b = 21, C = 94° choices: c ≈ 24.2, A ≈ 32.7°, B ≈ 53.3° c ≈ 24.9, A ≈ 54.3°, B ≈ 31.7° c ≈ 24.2, A ≈ 28.7°, B ≈ 57.3° c ≈ 24.9, A ≈ 28.7°, B ≈ 57.3°

OpenStudy (anonymous):

Do you know the law of cosines? You can try that to find the first value. After that, just use the law of sines to complete the rest.

OpenStudy (anonymous):

can you show me step by step?

OpenStudy (anonymous):

\[c^2=a^2+b^2-2abcos(C)\] you are given a, b, and C, solve for c

OpenStudy (anonymous):

im confused... @J_Liu

OpenStudy (anonymous):

Given the values of a, b and C, you can calculate \[c^2\] = value and take the square root.

OpenStudy (anonymous):

\[c^2 = 12^2 + 21^2 - 2(12)(21)\cos(94)\]